OpenStudy (anonymous):

Investigate the sequence \[x_{n+1}=\frac{ 1 }{ x _{n}}\]

OpenStudy (perl):

so multiply both sides by xn xn^2 + xn = 1 , xn^2 + xn -1 = 0 this is a quadratic in xn

OpenStudy (perl):

xn = [ sqrt(5 ) +- 1] / 2

OpenStudy (perl):

xn = [ -1 + sqrt(5)] /2 , [-1 - sqrt(5)] / 2 ,

OpenStudy (perl):

oh, woops, didnt see the xn+1 , but it comes out the same

OpenStudy (perl):

hey you changed your question

OpenStudy (perl):

this new sequence oscillates

OpenStudy (asnaseer):

lets say the sequence started at some value \(x_0=a\). now list what the rest of the terms in the sequence would be given this starting point.

OpenStudy (perl):

the sequence was , i thought x_n+1 + 1 = 1/x_n

OpenStudy (perl):

ok sorry i misread, my font no good

OpenStudy (perl):

the sequence will be a, 1/a, a, 1/a, ...

OpenStudy (anonymous):

sorry, I do not understand.

OpenStudy (anonymous):

\[x_{1}=a+1???\]

OpenStudy (asnaseer):

the "n" subscript represent the term in the sequence. so when n=0, we have \(x_0\) which I said lets say it has some value "a": \(x_0=a\) so the next term will be \(x_1\) and we know we can work this out using your formula:\[x_{n+1}=\frac{1}{x_n}\]therefore:\[x_1=\frac{1}{x_0}=\frac{1}{a}\]does that make sense?

OpenStudy (asnaseer):

the formula you have is basically saying: next term = 1 divided by previous term

OpenStudy (anonymous):

@asnaseer I kept getting 1/a

OpenStudy (asnaseer):

lets see what the next term would be...\[x_2=\frac{1}{x_1}=?\]

OpenStudy (anonymous):

x\[x{_{2}}=1/x _{1}=1/x _{0}\]

OpenStudy (asnaseer):

no

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

the formula says:\[x_{n+1}=\frac{1}{x_n}\]so, for n=0, we get:\[x_1=\frac{1}{x_0}\]for n=1, we get:\[x_2=\frac{1}{x_1}\]for n=2, we get:\[x_3=\frac{1}{x_2}\]etc...

OpenStudy (asnaseer):

and we started with: \(x_0=a\). so: \(\displaystyle x_1=\frac{1}{x_0}=\frac{1}{a}\) and: \(\displaystyle x_2=\frac{1}{x_1}=\frac{1}{\frac{1}{a}}=1\div\frac{1}{a}=1\times\frac{a}{1}=a\)

OpenStudy (anonymous):

yeah I get it.

OpenStudy (asnaseer):

so the first three terms of the sequence would be:\[a,\frac{1}{a},a\]

OpenStudy (asnaseer):

hopefully you can now see that it just oscillates between \(a\) and \(\displaystyle\frac{1}{a}\)

OpenStudy (asnaseer):

alos note that the sequence can NEVER have \(a=0\) otherwise you would have \(\displaystyle\frac{1}{0}=\infty\)

OpenStudy (anonymous):

yes I see that.

OpenStudy (asnaseer):

another interesting observation you could make is what happens if \(a=1\) or \(a=-1\)?

OpenStudy (anonymous):

I will try that now.

OpenStudy (anonymous):

is that \[x_{1}=a??\]

OpenStudy (asnaseer):

no, if \(x_0=a\) then \(\displaystyle x_1=\frac{1}{a}\)

OpenStudy (asnaseer):

so, if a=1, what will \(x_0=?\)

OpenStudy (asnaseer):

and what will \(x_1=?\)

OpenStudy (anonymous):

confused.

OpenStudy (asnaseer):

remember your formula generates the following sequence if \(x_0=a\):\[a,\frac{1}{a},a,\frac{1}{a},a,\frac{1}{a},...\]do you agree with this?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

so, if say we let \(a=2\), then we get the sequence:\[2,\frac{1}{2},2,\frac{1}{2},2,\frac{1}{2},...\]agreed?

OpenStudy (anonymous):

so \[x _{0}=1\]

OpenStudy (anonymous):

and \[x_{1}=1/1=1?\] if a=1?

OpenStudy (asnaseer):

yes, so what will the sequence look like if \(a=1\)?

OpenStudy (anonymous):

111111111

OpenStudy (asnaseer):

remember to put comma's between each term

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

so if \(a=1\) we get a sequence were every term equals 1. similarly, if \(a=-1\), we get a sequence where every term is equal to -1.

OpenStudy (anonymous):

thanks you. very much,

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

very interesting math.

OpenStudy (asnaseer):

I'm glad you find it interesting - it makes the subject much easier to learn if you find it interesting. :)